Let X be the random variable such that : P(X = –2) = P(X = –1), P(X = 2) = P(X = 1) and P(X > 0) = P(X < 0) = P(X = 0). Obtain the probability mass function of X and its distribution functions


1
Expert's answer
2022-03-02T03:23:13-0500
"P(X<0)=P(X = \u20132) +P(X = \u20131)"

"P(X>0)=P(X = 2) +P(X = 1)"

"P(X<0)+P(X = 0) +P(X>0)=1"

Then


"P(X<0)=P(X = 0) =P(X>0)=1\/3"

"P(X=-2)=P(X = -1)=1\/6"

"P(X=2)=P(X = 1)=1\/6"

Probability mass function


"p(x) = \begin{cases}
 1/6 &\text{if } x=-2 \\
 1/6 &\text{if } x=-1 \\
 1/3 &\text{if } x=0 \\
 1/6 &\text{if } x=1 \\
 1/6 &\text{if } x=2 \\
\end{cases}"

The cumulative distribution function



"F(x) = \begin{cases}
 0 & x<-2 \\
 1/6 & -2\le x<-1 \\
 1/3 &-1\le x<0 \\
 2/3 & 0\le x<1 \\
 5/6 & 1\le x<2 \\
 1 & x\ge2 \\
\end{cases}"